# MOLA: Enhancing Industrial Process Monitoring Using Multi-Block Orthogonal Long Short-Term Memory Autoencoder

### v1

@@ 3 The MOLA Fault Dection Framework
Now that all components of our proposed process monitoring framework are introduced, we now move on to discussing how these components are integrated in our MOLA framework during offline learning and online monitoring stages. The overall flowchart of MOLA is shown in Figure 4.
@@ 2. Step 2:
    Following the block assignments, standardized online data are sent to their correspoding block’s OLAE model to obtain the codes and monitoring statistics $T^{2}$ and $W$.
@@ 4. Step 4:
    Using the adaptive W-BF method again, we obtain the process-level PFI from local statistics of all blocks. If the PFI is greater than our pre-defined significance level $\alpha$ (0.01), a fault is declare; otherwise, the process is under normal operation and the monitoring continues.
@@ 4 Case Study
The Tennessee Eastman Process (TEP) is a simulation process developed by the Eastman Chemical Company based on an actual chemical process [37]. The TEP has been widely adopted as a benchmark for chemical process control, optimization, and monitoring. As illustrated in the process flow diagram of Figure 5, the TEP contains five major unit operations, which are associated with 12 manipulated variables and 41 measured variables in total. Among them, 31 variables (listed in Table 1) are typically selected to conduct process monitoring, as the remaining variables have relatively low sampling frequencies. Leveraging the topological structure and process knowledge of the TEP process, we partition the entire process into four distinct blocks ($N=4$) described in Table 2. In this work, the simulation data are generated from a Simulink implementation with a sampling frequency of one minute [38]. Initially, a dataset comprising 60,000 samples under normal operating conditions is generated and used for offline learning. Subsequently, additional 500 datasets, each containing 2,400 samples, are generated to determine the control limits in process monitoring models. Furthermore, this Simulink implementation can simulate 20 types of process faults (see Table 3) as test datasets for process online monitoring. It is worth noting that any process fault is introduced at the 600th sample in each test dataset.
@@ 20 Unknown Unknown
We compare our proposed framework with other process monitoring models, including PCA, AE, LSTM AE, Block PCA, and Block LSTM AE. Tables 4 and 5 show the performance of these models in terms of Fault Detection Delay (FDD) and Fault Detection Rate (FDR), respectively. It can be seen that, in general, the incorporation of a multi-block monitoring strategy not only reduces the FDD, but also increases the FDR of original process monitoring methods. This validates the effectiveness of our proposed multi-block method and data fusion technique. In particular, our MOLA framework exhibits superior performance over all other frameworks in terms of FDR and FDD. For the vast majority of fault cases, particularly fault 13, 15, 16, 24, and 18, MOLA can detect faults tens to hundreds of minutes ahead of other methods, providing invaluable time buffers for engineers and operators to take appropriate control actions. For a few faults, such as faults 3, 11, 14, and 20, MOLA falls behind the best-performing method by just a few minutes. However, considering that the differences are small and that MOLA achieves significant improvements in FDR in these faults, MOLA still outperforms other methods by a considerable margin. For example, for faults 3, 5, 9, 15, and 16, while other methods perform poorly on these faults, MOLA increases the FRD by 4.6 to 30 times. Note that faults 3, 9, and 15 are known to be notoriously difficult to detect due to their intricate process dynamics. Meanwhile, the FDR for MOLA on these three faults are between 88.3 and 98.5%, demonstrating MOLA’s exceptional capabilities in tackling challenging fault detection scenarios.
@@
We now take a close look at fault 3, whose process monitoring results for various methods are summarized in Figure 6. As we can see, while the LSTM AE-based approach successfully detects the fault at the 627th sample (i.e., 27 samples after the process fault is introduced into the simulation), its monitoring statistic only briefly exceeds the control limit and fails to raise a continuous alarm. Similarly, most other methods are unable to raise any valid alarm for this fault, as in practice alarms would only be considered as effective when the monitoring statistic exceeds the control limit multiple consecutive times (e.g., three). If a process monitoring model fails to continuous report faults, plant operators may erroneously assume that the alarm is false or the process has returned back to normal. On the other hand, our MOLA framework captures the fault starting at the 630th sample, and subsequently raises alarms continuously, making it a useful process monitoring framework in practice.
@@
4.1 Ablation studies evaluating contribution of MOLA components to its process monitoring performance
@@
We first validate the effectiveness of MOLA in extracting non-redundant features. The Maximal Information Coefficient (MIC) [39], ranging from 0 to 1, serves as a quantitative measure of the correlation between two variables. An MIC value of 0 between two variables indicates their mutual independence, whereas as MIC value approaches 1 it suggests the presence of a strong correlation between the two variables. Here, we use MIC as an indicator to assess the correlation among the extracted features. Specifically, we calculate the MIC values between features extracted by both the LSTM AE and the MOLA and summarize the results in Table 7. It is clear that the MIC values among features extracted by MOLA are close to 0, whereas those among features extracted by LSTM AE are close to 1. Thus, MOLA demonstrates a significant advantage in extracting non-redundant features. This is reflected by the comparison of process monitoring performance at the block level for LSTM AE and MOLA, both of which use $T^{2}$ as the monitoring statistic. As shown in Table 8, MOLA is able to detect faults earlier than LSTM AE in nearly all blocks.
@@
No. Block 1 Block 2 Block 3 Block 4 LSTM AE MOLA LSTM AE MOLA LSTM AE MOLA LSTM AE MOLA MIC(F1,F2) 0.2637 0.0337 0.0555 0.0541 0.1278 0.0706 0.9752 0.0405 MIC(F1,F3) 0.7957 0.1558 0.0708 0.0754 0.0427 0.0590 0.6531 0.0586 MIC(F1,F4) 0.3043 0.0628 0.1924 0.0908 0.0901 0.0867 0.7901 0.0547 MIC(F1,F5) 0.7275 0.0256 0.0774 0.1182 0.1287 0.0978 0.9610 0.0438 MIC(F2,F3) 0.7238 0.0723 0.4757 0.0879 0.0712 0.0938 0.9175 0.0789 MIC(F2,F4) 0.6029 0.1292 0.2940 0.1075 0.0663 0.0916 0.8663 0.0664 MIC(F2,F5) 0.5836 0.0946 0.5009 0.0918 0.1219 0.0559 0.9773 0.0550 MIC(F3,F4) 0.5106 0.1165 0.3697 0.0595 0.0352 0.0347 0.8658 0.0480 MIC(F3,F5) 0.9525 0.0948 0.9523 0.0742 0.0523 0.0387 0.8695 0.0417 MIC(F4,F5) 0.4059 0.0397 0.3442 0.0442 0.0944 0.1517 0.5427 0.0598
*F1, F2, F3, F4 and F5 represent the features extracted from the latent space.
@@
Fault No. Block 1 Block 2 Block 3 Block 4 LSTM AE MOLA LSTM AE MOLA LSTM AE MOLA LSTM AE MOLA 1 21 18 4 1 4 3 3 2 2 117 118 21 17 22 16 18 9 3 139 573 573 4 0 0 10 10 44 5 1 3 2 3 6 0 0 1 0 3 3 2 2 7 0 0 1 1 0 0 8 56 36 182 28 32 32 27 24 9 23 233 61 61 62 10 380 170 170 74 70 11 36 36 47 41 63 49 12 1553 85 60 48 60 65 65 13 249 245 97 74 89 84 106 9 14 202 2 2 56 26 15 82 192 191 209 16 54 18 19 17 56 54 118 117 140 57 18 345 345 263 260 346 346 19 409 244 25 20 12 11 20 164 162 212 177 129 129 123 124
@@
Next, we investigate the benefits of introducing the quantile-based CUSUM method and adaptive W-BF strategy on process monitoring performance. By designing ablation experiments, we quantify the enhancements and present the results in Table 9. Specifically, “Full MOLA” represents the complete MOLA framework presented earlier, “MOLA no CUSUM” means the MOLA framework without implementing the CUSUM procedure, and “MOLA no BF” means the MOLA framework without implementing adaptive W-BF strategy. Table 9 shows that introducing adaptive W-BF significantly improves the fault detection speed for faults 3, 15, and 18, and improves fault detection rate for all faults other than 1, 4, 6, 7, 10, and 14. For instance, for fault 3, the weights assigned by adaptive W-BF strategy at the 630th sample (i.e., 30 samples after fault is introduced) are 0.2478, 0.4531, 0.1432, and 0.1559 for blocks 1 through 4, respectively. Block 2, which has the highest weight to PFI among all blocks, also turns out to be the first block where an alarm is raised. Thus, the adaptive W-BF method automatically prompts the FPI to give higher attention to blocks that detect anomalies early, which helps enhancing the the fault detection rate and speed.
@@
20 124 0.7933 124 0.7917 124 0.7933
Meanwhile, a comparison between “full MOLA” and “MOLA no CUSUM” shows that incorporating quantile-based CUSUM procedure greatly enhances the fault detection rate and speed for hard-to-detect faults of 3, 9, and 15. Again, taking fault 3 as an example. As shown in Figure 8, under normal and faulty conditions, the numerical range of one of the specific feature extracted from block 2 remains largely the same, making traditional distance-based monitoring statistics such as $T^{2}$ struggle to detect the fault. However, when plotting the distribution of the features under normal and faulty conditions, considerable changes can be noticed. Therefore, quantile-based multivariate CUSUM method, which directly targets the detection of distribution changes, significantly improves the accuracy and speed of fault detection under these scenarios. Furthermore, as shown in Figure 9, quantile-based CUSUM statistic can consistently detect fault 3 by continuously accumulating the monitoring statistic, which offers an additional advantage over other monitoring methods.
@@ 5 Conclusions
In this work, we develop a novel process monitoring framework named MOLA for large-scale industrial processes. By adopting a multi-block monitoring strategy, MOLA successfully addresses the challenges posed by the scale and complexity of modern industrial processes. MOLA can extract dynamic orthogonal latent features, making sure that the most essential, non-redundant features are identified and extracted. In addition, MOLA incorporates a quantile-based multivariate CUSUM method which enhances the ability to detect faults characterized by subtle changes in feature distributions. Furthermore, the adaptive weight-based Bayesian fusion strategy enhances fault detection rate and speed. We remark that these methods lead to synergistic improvement in process monitoring performance when they are integrated together in the MOLA framework. Case study results on the TEP problem indicate that MOLA not only significantly improves fault detection rates and speeds but also successfully detects faults that are considered to be difficult to detect in prior research, thereby opening up many exciting opportunities for fast, accurate, and reliable industrial process monitoring applications.
Despite these achievements, we remark that there is still room for improvement for our MOLA framework. For instance, the current block assignment technique is solely based on our expert knowledge without fully considering the correlations among blocks, which may have an impact on both local and global monitoring performance. Therefore, future research will focus on exploring more scientific and reasonable methods for dividing process sub-blocks, taking into account the correlations between sub-blocks to further optimize the process monitoring model. This endeavor aims to further enhance its monitoring performance and robustness in complex industrial processes.

### v2

@@ 3 The MOLA Fault Dection Framework
Now that all components of our proposed process monitoring framework have been introduced, we will move on to discuss how these components are integrated with our MOLA framework during offline learning and online monitoring stages. The overall flowchart of MOLA is shown in Figure 4.
@@ 2. Step 2:
    Following the block assignments, standardized online data are sent to their corresponding block’s OLAE model to obtain the codes and monitoring statistics $T^{2}$ and $W$.
@@ 4. Step 4:
    Using the adaptive W-BF method again, we obtain the process-level PFI from the local statistics of all blocks. If the PFI is greater than our pre-defined significance level $\alpha$ (0.01), a fault is declared; otherwise, the process is under normal operation and the monitoring continues.
@@ 4 Case Study
The Tennessee Eastman Process (TEP) is a simulation process developed by the Eastman Chemical Company based on an actual chemical process [37]. The TEP has been widely adopted as a benchmark for chemical process control, optimization, and monitoring. As illustrated in the process flow diagram of Figure 5, the TEP contains five major unit operations, which are associated with 12 manipulated variables and 41 measured variables in total. Among them, 31 variables (listed in Table 2) are typically selected to conduct process monitoring, as the remaining variables have relatively low sampling frequencies. In multi-block process monitoring methods, the choice of block division and assignment approach directly influences the process monitoring performance. Existing block division methods include process knowledge-based methods, variable relation-based methods, and fault information-aided methods [38]. In this work, we adopt the process knowledge-based approach developed by Zhu et al. [39] and divide the entire process into four distinct blocks ($N=4$) as outlined in Table 3. The basic idea is to assign process variables associated with the same equipment into a single block. And due to the relatively small number of process variables being measured for the condenser and compressor, we assign these variables to the nearest separator block.
In this work, the simulation data are generated from a Simulink implementation with a sampling frequency of one minute [40]. Initially, a dataset comprising 60,000 samples under normal operating conditions is generated and used for offline learning. Subsequently, an additional 500 datasets, each containing 2,400 samples, are generated to determine the control limits in process monitoring models. Furthermore, this Simulink implementation can simulate 20 types of process faults (see Table 4) as test datasets for process online monitoring. It is worth noting that any process fault is introduced at the 600th sample in each test dataset.
@@ 20 Unknown Unknown
We compare our proposed framework with other process monitoring models, including PCA, AE, LSTM AE, Block PCA, and Block LSTM AE. The complete structure of each neural network implemented in our framework is listed in Table 5. In this work, all neural networks use the Rectified Linear Unit (ReLU) activation function. The number of iterations is determined using early stopping criteria. Specifically, we stop the training process when the loss function value on the validation dataset does not decrease for consecutive 30 iterations. The detailed structure for different neural networks is illustrated in Table 5. Tables 6 and 7 show the performance of these models in terms of Fault Detection Delay (FDD) and Fault Detection Rate (FDR), respectively. It can be seen that, in general, the incorporation of a multi-block monitoring strategy not only reduces the FDD but also increases the FDR of original process monitoring methods. This validates the effectiveness of our proposed multi-block method and data fusion technique. In particular, our MOLA framework exhibits superior performance over all other frameworks in terms of FDR and FDD. For the vast majority of fault cases, particularly faults 13, 15, 16, 24, and 18, MOLA can detect faults tens to hundreds of minutes ahead of other methods, providing invaluable time buffers for engineers and operators to take appropriate control actions. For a few faults, such as faults 3, 11, 14, and 20, MOLA falls behind the best-performing method by just a few minutes. However, considering that the differences are small and that MOLA achieves significant improvements in FDR in these faults, MOLA still outperforms other methods by a considerable margin. For example, for faults 3, 5, 9, 15, and 16, while other methods perform poorly on these faults, MOLA increases the FRD by 4.6 to 30 times. Note that faults 3, 9, and 15 are known to be notoriously difficult to detect due to their intricate process dynamics. Meanwhile, the fault detection rates for MOLA on these three faults range between 88.3 and 98.5%, demonstrating MOLA’s exceptional capabilities in tackling challenging fault detection scenarios.
Table 5: Detailed structure for different neural networks.
Model Structure Activation function
AE FC(64)-FC(16)-FC(64) ReLU
LSTM AE LSTM(64)-FC(16)-LSTM(64) ReLU
Block LSTM AE LSTM(15)-FC(5)-LSTM(15) ReLU
OLAE LSTM(15)-FC(5)-LSTM(15) ReLU
@@
We now take a close look at fault 3, whose process monitoring results for various methods are summarized in Figure 6. As we can see, while the LSTM AE-based approach successfully detects the fault at the 627th sample (i.e., 27 samples after the process fault is introduced into the simulation), its monitoring statistic only briefly exceeds the control limit and fails to raise a continuous alarm. Similarly, most other methods are unable to raise any valid alarm for this fault, as in practice alarms would only be considered effective when the monitoring statistic exceeds the control limit multiple consecutive times (e.g., three)[41]. If a process monitoring model fails to continuously report faults, plant operators may erroneously assume that the alarm is false or the process has returned back to normal. On the other hand, our MOLA framework captures the fault starting at the 630th sample, and subsequently raises alarms continuously, making it a useful process monitoring framework in practice.
@@
4.1 Ablation studies evaluating the contribution of MOLA components to its process monitoring performance
@@
We first validate the effectiveness of MOLA in extracting non-redundant features. The Maximal Information Coefficient (MIC) [42], ranging from 0 to 1, serves as a quantitative measure of the correlation between two variables. A MIC value of 0 between two variables indicates their mutual independence, whereas as MIC value approaches 1 it suggests the presence of a strong correlation between the two variables. Here, we use MIC as an indicator to assess the correlation among the extracted features. Specifically, we calculate the MIC values between any two features extracted by both the LSTM AE and the MOLA and summarize the results in Table 9. Clearly, the MIC values for features extracted by MOLA are close to 0, whereas the MIC values for features extracted by LSTM AE are significantly larger. This indicates that MOLA demonstrates a significant advantage in extracting non-redundant features. As illustrated in Table 10, this leads to faster fault detection speed at the block level for MOLA compared to LSTM AE, when both methods use $T^{2}$ as the monitoring statistic. Meanwhile, MOLA without W-BF and CUSUM can detect process faults earlier than LSTM AE. This result indicates that introducing orthogonality constraints can facilitate faster detection of process faults. This is primarily because the redundancies among features extracted by traditional autoencoders may obscure the underlying data structure, making faults harder to detect. Instead, our proposed OLAE only extracts orthogonal features, making fault information more prominent and thus easier to detect.
@@
No. Block 1 Block 2 Block 3 Block 4 Block LSTM AE MOLA w/o BF and CUSUM LSTM AE OLAE LSTM AE OLAE LSTM AE OLAE LSTM AE OLAE 1 21 18 4 1 4 3 3 2 4 2 2 117 118 21 17 22 16 18 9 18 16 3 - - - 139 573 573 - - 572 160 4 - - 0 0 10 10 - 44 0 0 5 - - - 1 3 2 - 3 3 2 6 0 0 1 0 3 3 2 2 0 0 7 - - 0 0 1 1 0 0 0 0 8 56 36 182 28 32 32 27 24 27 27 9 - 23 233 61 - 61 - 62 233 61 10 - - - 380 170 170 74 70 74 70 11 - - 36 36 47 41 63 49 36 36 12 - 553 85 60 48 60 65 65 48 60 13 249 245 97 74 89 84 106 9 91 84 14 - 202 2 2 - 56 - 26 2 2 15 - - - 82 192 191 - 209 191 192 16 - - - - 54 18 - 19 - 25 17 - - 56 54 118 117 140 57 57 56 18 - - 345 345 263 260 346 346 268 260 19 - - 409 244 25 20 12 11 12 12 20 164 162 212 177 129 129 123 124 123 127
@@
Fault Block 1 Block 2 Block 3 Block 4 No. LSTM AE MOLA LSTM AE MOLA LSTM AE MOLA LSTM AE MOLA 1 21 18 4 1 4 3 3 2 2 117 118 21 17 22 16 18 9 3 139 573 573 4 0 0 10 10 44 5 1 3 2 3 6 0 0 1 0 3 3 2 2 7 0 0 1 1 0 0 8 56 36 182 28 32 32 27 24 9 23 233 61 61 62 10 380 170 170 74 70 11 36 36 47 41 63 49 12 1553 85 60 48 60 65 65 13 249 245 97 74 89 84 106 9 14 202 2 2 56 26 15 82 192 191 209 16 54 18 19 17 56 54 118 117 140 57 18 345 345 263 260 346 346 19 409 244 25 20 12 11 20 164 162 212 177 129 129 123 124
@@
Next, we investigate the benefits of introducing the quantile-based CUSUM method and adaptive W-BF strategy on process monitoring performance. By designing ablation experiments, we quantify the enhancements and present the results in Table 11. Specifically, “Full MOLA” represents the complete MOLA framework presented earlier, “MOLA no CUSUM” means the MOLA framework without implementing the CUSUM procedure, and “MOLA no BF” means the MOLA framework without implementing adaptive W-BF strategy. Table 11 shows that introducing adaptive W-BF significantly improves the fault detection speed for faults 13 and 15 and enhances fault detection rate for all faults other than 1, 4, 6, 7, 10, and 14. For instance, for fault 13, the weights assigned to blocks 1 through 4 by adaptive W-BF strategy at the 609th time step (i.e., 9 samples after fault is introduced) are 0.2396, 0.1372, 0.1542, and 0.4689, respectively. Block 4, which has the highest weight to PFI among all blocks, also turns out to be the first block where an alarm is raised. Thus, the adaptive W-BF method automatically prompts the FPI to give higher attention to blocks that detect anomalies earlier, which enhances both fault detection rate and speed.
We remark that the W-BF method is particularly attractive in monitoring modern, large-scale industrial processes, whose scale and complexity require process monitoring to be performed in a distributed manner. Since process faults typically occurs locally, by decomposing the overall process into multiple blocks, our W-BF method ensures that, as a fault occurs, process monitoring performance is sensitive to the faulty blocks and will not be negatively impacted by the surveillance of non-faulty blocks.
@@
20 124 124 124 0.7917 124 0.7933
Average 30.45 0.8734 30.85 0.8658 56.25 0.6855
Meanwhile, a comparison between “full MOLA” and “MOLA no CUSUM” shows that incorporating quantile-based CUSUM procedure greatly enhances the fault detection rate and speed for hard-to-detect faults of 3, 9, and 15. Again, take fault 3 as an example. As shown in Figure 8, under normal and faulty conditions, the numerical range of one of the specific features extracted from block 2 remains largely the same, making traditional distance-based monitoring statistics such as $T^{2}$ struggle to detect the fault. However, when plotting the distribution of the features under normal and faulty conditions, considerable changes are noticed. Therefore, the quantile-based multivariate CUSUM method, which directly targets the detection of distribution changes, significantly improves the accuracy and speed of fault detection under these scenarios. Furthermore, as shown in Figure 11, the quantile-based CUSUM statistic continuously accumulates as fault 3 occurs and will not go below the control limit once an alarm is raised. This causes less confusion among plant operators in interpreting alarms and thus offers an additional advantage over other monitoring methods.
@@ 5 Conclusions
In this work, we develop a novel process monitoring framework named MOLA for large-scale industrial processes. By adopting a multi-block monitoring strategy, MOLA successfully addresses the challenges posed by the scale and complexity of modern industrial processes. MOLA can extract dynamic orthogonal latent features, making sure that the most essential, non-redundant features are identified and extracted. In addition, MOLA incorporates a quantile-based multivariate CUSUM method which enhances the ability to detect faults characterized by subtle changes in feature distributions. Furthermore, the adaptive weight-based Bayesian fusion strategy enhances fault detection rate and speed. We remark that these methods lead to synergistic improvement in process monitoring performance when they are integrated with the MOLA framework. Case study results on the TEP problem indicate that MOLA not only significantly improves fault detection rates and speeds but also successfully detects faults that are considered difficult to be detected in prior research, thereby opening up many exciting opportunities for fast, accurate, and reliable industrial process monitoring applications.
Despite these achievements, we remark that there is still room for improvement for our MOLA framework. For instance, the current block assignment technique is solely based on our expert knowledge without fully considering the correlations among blocks, which may have an impact on both local and global monitoring performance. Therefore, future research will focus on exploring more systematic methods for dividing process sub-blocks, taking into account the correlations between sub-blocks to further optimize the process monitoring model. For example, we plan to perform correlation analysis on process variables prior to assigning them to different blocks. This way, the dynamic features extracted not only capture the characteristics of process data within individual blocks, but also incorporate the interconnections across different blocks. This will further enhance its monitoring performance and robustness in complex industrial processes.